OpenStudy (anonymous):

Help please........ Point O is the center of the circle. Which is greater or are they equal? A. The ratio of the length of minor arc AB to major arc AB B. 1/6

OpenStudy (anonymous):

NOPE

OpenStudy (anonymous):



OpenStudy (anonymous):

nope its B

OpenStudy (anonymous):

I need to know why i don't need the answer.

OpenStudy (anonymous):

oh well my explaination is that it is 1/6 because it makes sense

OpenStudy (anonymous):

@Nnesha

OpenStudy (ytrewqmiswi):

let the radius be r so circumference=2(pi)r let the length of major arc be x so length of minor arc be 2(pi)r-x ratio of lengths= [2(pi)r-x]/x =2(pi)r /x -1 and x =2(pi)r x 60/360 =2(pi)r/6 so ratio is -->2(pi)r/[2(pi)r/6] -1 =6-1 =5 so 5>1/6 a is bigger